# Dynamically Adaptive Reasoning via LLM-Guided MCTS for Efficient and Context-Aware KGQA

### v3

@@ Introduction
Performance Comparison
We report the experimental results of DAMR in Table 1, benchmarking its performance against state-of-the-art baselines across KGQA datasets. From the results, we find that: (1) Semantic parsing and retrieval-based methods serve as early foundations for KGQA by extracting subgraphs and capturing structural semantics. However, embedding-based models struggle with complex relational patterns, while retrieval-based methods rely on rigid pipelines that limit generalization. In contrast, LLM with KG approaches combine the language understanding of LLMs with structured reasoning over KGs, enabling more flexible path exploration and improved adaptability to diverse, multi-hop queries. (2) General-purpose LLMs, such as ChatGPT and Alpaca-7B, show basic reasoning ability but often perform worse than methods that combine LLMs with KGs in KGQA tasks. This is mainly because they are not grounded in domain-specific knowledge, making them more likely to produce incorrect or made-up answers. (3) DAMR consistently outperforms all baselines across both datasets, showcasing its strong reasoning capability. This superior performance is driven by its integration of an LLM-based planner, which selectively retrieves relevant relations to reduce noise and guide the search toward high-quality reasoning paths, and a path evaluation model that is dynamically fine-tuned during search to capture semantic differences among candidate paths and accurately rank those most likely to yield correct answers.
Efficiency Analysis
Ablation Study
We conduct ablation studies to examine the key components in DAMR: (1) DAMR w/o PE: It removes the path evaluation module; (2) DAMR w/o FT: disables the fine-tuning mechanism for the path evaluation module; (3) DAMR w/ GPT 4.1: replaces the context-aware path evaluation module with a general LLM.
Experimental results are summarized in Table 3. From the results, we find that: (1) Removing the path evaluation module (DAMR w/o PE) leads to a noticeable performance drop on both datasets, highlighting its critical role in guiding the search process. Without this component, the model cannot effectively assess or rank candidate paths, leading to suboptimal reasoning and degraded answer accuracy. (2) Compared to DAMR w/o FT, the proposed DAMR consistently achieves superior results on both datasets, highlighting the importance of the finetuning mechanism in the path evaluation module. This mechanism enables the model to adapt to the evolving distribution of explored paths, improving its ability to distinguish between plausible and implausible reasoning trajectories. (3) Replacing the context-aware path evaluation module with general LLMs leads to degraded performance, confirming the advantage of our fine-tuned path scorer. By capturing fine-grained semantic distinctions among candidate paths, it provides more accurate evaluation signals, thereby enhancing the overall search effectiveness.
@@
Table 4: Performance of DAMR using different LLM-based planners as backbones on the WebQSP and CWQ datasets. Bold values denote the best results.
@@
Table 5: Case study of DAMR. We highlight the correct answers in Bold and the wrong answers in underline.
Question What structure opened in 1922 in Hollywood, CA?
Answer Grauman’s Egyptian Theatre
Llama-2-13B The Grauman’s Chinese Theatre, a historic movie palace, opened on November 18, 1922, in Hollywood, CA. It was built by Sid Grauman and Charles E. Toberman and has since become a famous landmark and tourist attraction.
Qwen-3-14B The TCL Chinese Theatre, originally known as the Chinese Grand Theatre, opened in 1922 in Hollywood, CA. It is a historic movie theater renowned for its unique architecture and celebrity handprints and footprints.
GPT 4.1 The Hollywood Bowl, an iconic amphitheater in Hollywood, California, opened in 1922. It is renowned for its distinctive band shell and has hosted numerous concerts and events, becoming a significant cultural landmark in the area.
GPT 4.1-mini The Hollywood Bowl, an iconic amphitheater in Hollywood, California, opened in 1922 and has since been a renowned venue for music performances and cultural events.
DAMR Path 1: Entity (id: 83076) $\rightarrow$ location.location.events $\rightarrow$ time.event.locations $\rightarrow$ travel.travel_destination.tourist_attractions $\rightarrow$ Grauman’s Egyptian Theatre. Path 2: Entity (id: 83076) $\rightarrow$ travel.travel_destination.tourist_attractions $\rightarrow$ Grauman’s Egyptian Theatre.
Sensitivity Analysis
Figure 2 illustrates how $k$ and $L$ affect the performance of DAMR on the WebQSP and CWQ datasets. We vary $k$ and $L$ within the range of $\{2,3,4,5\}$. From the results, we observe that: (1) As shown in Figure 2(a), increasing $k$ initially leads to performance gains, which then stabilize before experiencing a slight decline. While larger $k$ values encourage broader relational exploration, they may also introduce irrelevant candidates and increased computational cost. Conversely, smaller $k$ restrict the diversity of the search. To balance these trade-offs, we select a moderate $k=3$ as the default setting. (2) As shown in Figure 2(b), on the WebQSP dataset, performance improves from $L=2$ to $3$, then fluctuates between $L=3$ and $5$, suggesting limited gains beyond three hops. In contrast, performance on the CWQ dataset steadily increases up to $L=4$ before slightly declining at $L=5$, reflecting its need for deeper reasoning due to more complex questions. Balancing effectiveness and efficiency across both datasets, we set $L=4$ as the default path length in all experiments. More results are provided in Appendix E.
Impact of Different LLMs
@@
To evaluate the impact of different LLM-based planners within the DAMR framework, we compare several backbones including Llama2 13B (Roque 2025), Qwen3 14B (Team 2024), GPT 4.1 mini, and GPT 4.1, as shown in Table 4. Across both datasets, stronger LLMs consistently yield higher F1 and Hits scores, with GPT 4.1 achieving the best performance on all metrics. This highlights the critical role of advanced LLMs in guiding relation selection and reasoning path expansion. The results confirm that improved language modeling and semantic understanding capabilities directly enhance KGQA accuracy. Overall, these findings emphasize the importance of backbone selection and further validate the design of DAMR, which leverages powerful LLMs for robust and effective multi-hop reasoning.
Case study
Table 5 presents a case study comparing the reasoning process of DAMR with four general LLMs: Llama-2-13B, Qwen-3-14B, GPT 4.1-mini, and GPT 4.1. While all baseline LLMs fail to identify the correct structure that opened in Hollywood in 1922, the proposed DAMR accurately finds Grauman’s Egyptian Theatre by explicitly traversing relation paths in the knowledge graph from two different reasoning paths. This example demonstrates that, although LLMs appear capable of answering the question, their responses can still be factually incorrect due to a lack of grounded knowledge. In contrast, DAMR consistently produces accurate and faithful answers by grounding its reasoning in KG and explicitly modeling reasoning paths. More studies can be found in Appendix E.
Conclusion
In this work, we present DAMR, a dynamically adaptive MCTS-based reasoning framework for complex KGQA tasks. DAMR incorporates an LLM-based planner to guide top-$k$ relation expansion, a context-aware path evaluator to assess reasoning paths without further LLM queries, and a dynamic refinement module that continually adapts the evaluator using pseudo-path supervision from MCTS rollouts. This modular design enables efficient yet accurate multi-hop reasoning by narrowing the search space, reducing redundant computation, and enhancing evaluation quality. Extensive experiments on WebQSP and CWQ confirm the effectiveness and efficiency of DAMR, making it a practical and scalable solution for real-world KGQA deployment.
Appendix A A. Algorithm
@@ Conclusion
Appendix B B. Complexity Analysis
The overall time complexity of the proposed DAMR framework is governed by its two primary online stages: the LLM-Guided MCTS Search and the interleaved Path-based Dynamic Refinement. In the MCTS Search phase, executed over $N$ iterations, the LLM-guided relation expansion incurs a total complexity of $\mathcal{O}\left(N\cdot T_{\text{LLM}}(k)\right)$, where $T_{\text{LLM}}(k)$ denotes the inference time of the LLM when provided with up to $k$ candidate relations from the Knowledge Graph (KG). The context-aware path evaluation performed during the simulation step introduces an additional cost of $\mathcal{O}\left(N\cdot k\cdot L^{3}\cdot d\right)$, where $L$ is the maximum path length and $d$ is the embedding dimension. In the Path-based Dynamic Refinement stage, the evaluator is fine-tuned for $N_{\text{FT}}$ steps, with a total complexity of $\mathcal{O}\left(N_{\text{FT}}\cdot L^{2}\cdot d\right)$. Consequently, the overall time complexity of DAMR is: $\mathcal{O}\left(N\cdot\left(T_{LLM}(k)+k\cdot L^{3}\cdot d\right)+N_{FT}\cdot L^{2}\cdot d\right)$.
Appendix C C. Datasets
Dataset Description
Table 6: Statistics of KGQA datasets.
@@
We conduct extensive experiments on two widely used multi-hop Knowledge Graph Question Answering (KGQA) benchmarks: WebQSP (Talmor and Berant 2018) and CWQ (Yih et al. 2016). The statistics of these two benchmarks can be found in Table 6, and their details are shown as follows:
  • The WebQuestionsSP (WebQSP) dataset is a widely adopted benchmark for evaluating single-hop and simple multi-hop KGQA (Yih et al. 2016). It consists of 4,837 natural language questions annotated with corresponding SPARQL queries over the Freebase knowledge graph. The dataset is partitioned into 2,848 training, 250 validation, and 1,639 test instances.
  • The ComplexWebQuestions (CWQ) dataset is a challenging benchmark designed for multi-hop KGQA (Talmor and Berant 2018). It comprises 34,689 questions derived from WebQuestionsSP, reformulated to include more complex and compositional queries. Each question typically requires multi-step reasoning over the Freebase knowledge graph, often involving conjunctions, comparatives, or nested logical structures. The dataset is divided into 27,639 training, 3,519 validation, and 3,531 test examples.
Data Processing
Following prior work (Shen et al. 2025; Long et al. 2025), we preprocess the datasets by constructing localized subgraphs centered around each question entity to reduce the size of the search space. Specifically, for each question in WebQSP (Yih et al. 2016) and CWQ (Talmor and Berant 2018), we extract a subgraph from the Freebase knowledge graph by including all triples within a predefined number of hops from the topic entity. This approach preserves the essential context required for multi-hop reasoning while significantly improving computational efficiency.
Appendix D D. Baselines
@@
    • KV-Mem: KV-Mem (Miller et al. 2016) introduce a neural architecture that stores facts as key-value pairs and enables question answering by attending over memory slots, directly retrieving relevant information to infer answers.
    • EmbedKGQA: EmbedKGQA (Saxena, Tripathi, and Talukdar 2020) enhances multi-hop question answering over knowledge graphs by leveraging pretrained knowledge base embeddings, enabling the model to reason over entity and relation representations without explicit path enumeration during answer prediction.
    • QGG: QGG (Lan and Jiang 2020) generates query graphs to answer multi-hop complex questions over knowledge bases, formulating question answering as query graph prediction and enabling structured reasoning through graph matching and path ranking mechanisms.
    • NSM: NSM (He et al. 2021) enhances multi-hop KBQA by leveraging intermediate supervision signals, decomposing questions into reasoning steps, and training a neural state machine to sequentially predict relations and entities for accurate path-based reasoning.
    • TransferNet: TransferNet (Shi et al. 2021) proposes a transparent framework for multi-hop QA over relational graphs by transferring question semantics to relation paths through interpretable path ranking and structured reasoning, enabling effective and explainable answer prediction.
    • KGT5: KGT5 (Saxena, Kochsiek, and Gemulla 2022) formulates knowledge graph completion and question answering as unified sequence-to-sequence tasks, leveraging pre-trained language models to jointly encode input queries and generate answer entities or triples in a flexible and end-to-end manner.
@@
    • GraftNet: GraftNet (Sun et al. 2018) proposes an early fusion framework that jointly encodes knowledge base facts and supporting text by constructing a heterogeneous graph, enabling effective reasoning through graph convolutional networks for open-domain question answering.
    • PullNet: PullNet (Sun, Bedrax-Weiss, and Cohen 2019) introduces an iterative retrieval mechanism that expands a query-specific subgraph by pulling relevant facts from both knowledge bases and text, enabling joint reasoning over heterogeneous evidence for open-domain question answering.
    • SR+NSM: SR+NSM (Zhang et al. 2022) enhances multi-hop KBQA by first retrieving a question-relevant subgraph and then performing symbolic reasoning over it using Neural Symbolic Machines, improving efficiency and accuracy through constrained and focused logical inference.
    • SR+NSM+E2E: SR+NSM+E2E (Zhang et al. 2022) extends SR+NSM by enabling end-to-end training that jointly optimizes subgraph retrieval and reasoning. This integration enhances model coherence and allows better alignment between retrieved subgraphs and final answer prediction.
@@
    • Flan-T5-xl: Flan-T5-xl (Chung et al. 2024) is an instruction-finetuned variant of the T5 model, trained on a diverse collection of tasks with natural language instructions. By leveraging large-scale instruction tuning, it improves zero-shot and few-shot performance across diverse NLP benchmarks.
    • Alpaca-7B: Alpaca-7B (Taori et al. 2023) is an instruction-following language model fine-tuned from LLaMA-7B using self-instruct techniques. It demonstrates strong zero-shot and few-shot performance by aligning with human instructions across various NLP tasks.
    • Llama3-8B: Llama3-8B (Dubey et al. 2024) is part of the LLaMA 3 family of models, designed for improved instruction following, reasoning, and code generation. Pretrained on a high-quality corpus and fine-tuned with supervised signals, it achieves strong performance across diverse benchmarks.
    • Qwen2.5-7B: Qwen2.5-7B (Team 2024) is a 7B-parameter instruction-tuned language model developed by Alibaba, optimized for tasks such as reasoning, code generation, and dialogue. It supports multi-turn conversation and demonstrates competitive performance on standard benchmarks.
    • ChatGPT: ChatGPT (Schulman et al. 2022) is a conversational AI developed by OpenAI, based on the GPT architecture. It is designed to understand natural language, engage in dialogue, answer questions, and assist with a wide range of tasks across domains.
    • ChatGPT+CoT: ChatGPT with Chain-of-Thought (CoT) (Wei et al. 2022) prompting enhances the model’s reasoning capabilities by encouraging it to generate intermediate reasoning steps before arriving at a final answer, improving performance on complex, multi-step problems.
@@
    • UniKGQA: UniKGQA (Jiang et al. 2022) is a unified framework that integrates retrieval and reasoning for multi-hop question answering over knowledge graphs, combining subgraph retrieval, query decomposition, and neural reasoning in an end-to-end manner.
@@
    • DECAF: DECAF (Yu et al. 2022) is a joint framework for question answering over knowledge bases that simultaneously decodes logical forms and answers. By leveraging dual supervision, it enhances both symbolic reasoning accuracy and direct answer prediction in a unified architecture.
@@
    • KD-CoT: KD-CoT (Wang et al. 2023) is a framework that enhances the faithfulness of large language models by guiding Chain-of-Thought reasoning with external knowledge, improving accuracy in knowledge-intensive question answering tasks.
@@
    • Nutrea: Nutrea (Choi et al. 2023) proposes a neural tree search framework for context-guided multi-hop KGQA. It incrementally constructs reasoning trees by integrating question semantics and graph context, enabling efficient exploration of multi-hop paths for accurate answer prediction.
@@
    • ToG: ToG (Sun et al. 2023) is a framework that enables large language models to perform deep and responsible reasoning over knowledge graphs by combining structured graph information with iterative thinking and verification mechanisms for reliable multi-hop QA.
@@
    • RoG: RoG (Luo et al. 2023) is a framework that enhances the faithfulness and interpretability of large language model reasoning by grounding multi-hop question answering on knowledge graphs, integrating symbolic path tracking with natural language generation.
@@
    • KAPING: KAPING (Baek, Aji, and Saffari 2023) introduces knowledge-augmented prompting by integrating structured triples into Chain-of-Thought (CoT) reasoning. It guides large language models to generate intermediate reasoning steps, enabling zero-shot multi-hop KGQA without task-specific fine-tuning.
@@
    • ReasoningLM: ReasoningLM (Jiang et al. 2023) enhances pre-trained language models for KGQA by injecting subgraph structures into the input representation. It enables structural reasoning over retrieved subgraphs through a reasoning-aware encoder, improving performance on complex multi-hop queries.
@@
    • FiDeLis: FiDeLis (Sui et al. 2024) proposes a faithfulness-aware KGQA framework that enhances reasoning consistency in LLMs by aligning generated logical forms with answer predictions. It introduces fidelity constraints to reduce hallucinations and improve answer correctness.
@@
    • GNN-RAG: GNN-RAG (Mavromatis and Karypis 2024) integrates graph neural networks with retrieval-augmented generation by encoding knowledge subgraphs into LLMs’ context. It enables structural reasoning over retrieved subgraphs, improving answer accuracy in KGQA through explicit graph-aware representations.
@@
    • DoG: DoG (Ma et al. 2025a) is a flexible and reliable reasoning framework that enables large language models to generate and evaluate multiple reasoning paths over knowledge graphs through a debate-style process, enhancing robustness and answer faithfulness.
@@
    • DuarL: DuarL (Liu et al. 2025) is a collaborative framework that integrates GNNs and LLMs for KGQA, where GNNs capture structural semantics and LLMs perform adaptive reasoning, enabling accurate and interpretable multi-hop QA.
@@
    • DP: DP (Ma et al. 2025b) is a trustworthy reasoning framework that guides large language models using prior knowledge from knowledge graphs. It iteratively verifies and refines reasoning paths to enhance faithfulness, robustness, and answer accuracy in KGQA.
@@
    • RwT: RwT (Shen et al. 2025) is a faithful KGQA framework that models multi-hop reasoning as tree-structured exploration over knowledge graphs, enabling large language models to generate interpretable reasoning paths and improve answer consistency and accuracy.
Question Where does greek language come from?
Answer Attic Group
Llama-2-13B The Greek language originates from the Proto-Indo-European language family, which spread to Greece around 2000 BCE. It evolved into the ancient Greek language, influenced by Mycenaean and Aeolic dialects, and eventually developed into Modern Greek.
Qwen-3-14B The Greek language originates from the Proto-Hellenic branch of the Indo-European language family. It evolved from the Mycenaean Greek of the second millennium BCE and became the foundation for Classical Greek, spoken in ancient Greece.
GPT 4.1 The Greek language originates from the Indo-European language family. It developed in the southern Balkans around the 3rd millennium BCE and is the oldest recorded living language, with written records dating back to the 15th century BCE (Linear B script).
GPT 4.1-mini The Greek language originates from the Hellenic branch of the Indo-European language family. It developed in the region of Greece and the Aegean around the 3rd millennium BCE, evolving through ancient, medieval, and modern stages.
DAMR Path 1: Entity (id: 120026) $\rightarrow$ base.rosetta.languoid.parent $\rightarrow$ Attic Group.
Table 7: Case study of DAMR. We highlight the correct answers in Bold and the wrong answers in underline.

### v4

@@ 4.4 Ablation Study
4.5 Sensitivity Analysis
Figure 2 illustrates how $k$ and $L$ affect the performance of DAMR on the WebQSP and CWQ datasets. We vary $k$ and $L$ within the range of $\{2,3,4,5\}$. From the results, we observe that: (1) As shown in Figure 2(a), increasing $k$ initially leads to performance gains, which then stabilize before experiencing a slight decline. While larger $k$ values encourage broader relational exploration, they may also introduce irrelevant candidates and increased computational cost. Conversely, smaller $k$ restrict the diversity of the search. To balance these trade-offs, we select a moderate $k=3$ as the default setting. (2) As shown in Figure 2(b), on the WebQSP dataset, performance improves from $L=2$ to $3$, then fluctuates between $L=3$ and $5$, suggesting limited gains beyond three hops. In contrast, performance on the CWQ dataset steadily increases up to $L=4$ before slightly declining at $L=5$, reflecting its need for deeper reasoning due to more complex questions. Balancing effectiveness and efficiency across both datasets, we set $L=4$ as the default path length in all experiments. More results are provided in Appendix  D.
4.6 Impact of Different LLMs
@@
Table 4: Performance of DAMR using different LLM-based planners as backbones on different datasets.
@@
To evaluate the impact of different LLM-based planners within the DAMR framework, we compare several backbones including Llama2-13B (roque2025evolution, ), Qwen3-14B (team2024qwen2, ), GPT 4.1 mini, and GPT 4.1, as shown in Table 4. Across both datasets, stronger LLMs consistently yield higher F1 and Hits scores, with GPT 4.1 achieving the best performance across all metrics. These results highlight the pivotal role of advanced LLMs in guiding relation selection and reasoning path expansion, where improved fluency, contextual awareness, and semantic precision translate into more accurate and faithful reasoning trajectories. Notably, the consistent performance gap between smaller and larger backbones illustrates the sensitivity of KGQA systems to the planner’s reasoning capability, reinforcing the necessity of leveraging high-capacity models when available. Overall, these findings emphasize the importance of backbone selection and further validate the design of DAMR, which capitalizes on powerful LLMs to achieve robust, generalizable, and effective multi-hop reasoning.
(a) One example from the WebQSP dataset.
(b) One example from the CWQ dataset.
Figure 3: Case studies of DAMR on the WebQSP and CWQ datasets. We highlight the correct answers in Red and the wrong answers in Blue.
4.7 Case study
Figure 3 presents detailed case studies on the WebQSP and CWQ datasets comparing the reasoning process of DAMR with four representative LLMs: Llama-2-13B, Qwen-3-14B, GPT 4.1-mini, and GPT 4.1. In Figure 3a, the baseline LLMs generate fluent and seemingly plausible answers such as “Proto-Indo-European” or “Proto-Hellenic” when asked about the origin of the Greek language, yet fail to identify the correct answer Attic Group. In contrast, DAMR accurately predicts the correct entity by explicitly traversing the base.rosetta.languoid.parent relation within the knowledge graph. In Figure 3b, none of the baseline models are able to identify the structure that opened in Hollywood in 1922, while the proposed DAMR successfully locates Grauman’s Egyptian Theatre by explicitly following relevant relation paths in the knowledge graph and integrating complementary reasoning trajectories. These examples collectively highlight the strength of DAMR in grounding its reasoning in the knowledge graph and explicitly modeling multi-step reasoning paths, enabling it to provide accurate and faithful answers to complex, ontology-specific questions that general-purpose LLMs often struggle to resolve.
5 Conclusion
In this work, we presented DAMR, a dynamically adaptive MCTS-based framework for knowledge graph question answering that integrates an LLM-guided planner, a context-aware path evaluator, and a dynamic pseudo-path refinement mechanism. By narrowing the search space, reducing redundant LLM calls, and continually adapting path evaluation, DAMR achieves efficient and accurate multi-hop reasoning. Extensive experiments on WebQSP and CWQ demonstrate that DAMR consistently surpasses state-of-the-art baselines in both performance and efficiency, while ablation and case studies highlight its interpretability and robustness. These results establish DAMR as a scalable solution for real-world KGQA and open avenues for extending adaptive symbolic–neural reasoning to broader domains such as scientific discovery and recommendation.
Appendix A Algorithm
@@
Appendix B Datasets
B.1 Dataset Description
@@
Table 5: Statistics of KGQA datasets.
We conduct extensive experiments on two widely used multi-hop Knowledge Graph Question Answering (KGQA) benchmarks: WebQSP [47] and CWQ [61]. The statistics of these two benchmarks can be found in Table 5, and their details are shown as follows:
  • The WebQuestionsSP (WebQSP) dataset is a widely adopted benchmark for evaluating single-hop and simple multi-hop KGQA [61]. It consists of 4,837 natural language questions annotated with corresponding SPARQL queries over the Freebase knowledge graph. The dataset is partitioned into 2,848 training, 250 validation, and 1,639 test instances.
  • The ComplexWebQuestions (CWQ) dataset is a challenging benchmark designed for multi-hop KGQA [47]. It comprises 34,689 questions derived from WebQuestionsSP, reformulated to include more complex and compositional queries. Each question typically requires multi-step reasoning over the Freebase knowledge graph, often involving conjunctions, comparatives, or nested logical structures. The dataset is divided into 27,639 training, 3,519 validation, and 3,531 test examples.
B.2 Data Processing
Following prior work [41, 26, 31], we preprocess the datasets by constructing localized subgraphs centered around each question entity to reduce the size of the search space. Specifically, for each question in WebQSP [61] and CWQ [47], we extract a subgraph from the Freebase knowledge graph by including all triples within a predefined number of hops from the topic entity. This approach preserves the essential context required for multi-hop reasoning while significantly improving computational efficiency.
Appendix C Baselines
@@
    • KV-Mem: KV-Mem [34] introduce a neural architecture that stores facts as key-value pairs and enables question answering by attending over memory slots, directly retrieving relevant information to infer answers.
    • EmbedKGQA: EmbedKGQA [39] enhances multi-hop question answering over knowledge graphs by leveraging pretrained knowledge base embeddings, enabling the model to reason over entity and relation representations without explicit path enumeration during answer prediction.
    • QGG: QGG [18] generates query graphs to answer multi-hop complex questions over knowledge bases, formulating question answering as query graph prediction and enabling structured reasoning through graph matching and path ranking mechanisms.
    • NSM: NSM [11] enhances multi-hop KBQA by leveraging intermediate supervision signals, decomposing questions into reasoning steps, and training a neural state machine to sequentially predict relations and entities for accurate path-based reasoning.
    • TransferNet: TransferNet [42] proposes a transparent framework for multi-hop QA over relational graphs by transferring question semantics to relation paths through interpretable path ranking and structured reasoning, enabling effective and explainable answer prediction.
    • KGT5: KGT5 [38] formulates knowledge graph completion and question answering as unified sequence-to-sequence tasks, leveraging pre-trained language models to jointly encode input queries and generate answer entities or triples in a flexible and end-to-end manner.
@@
    • GraftNet: GraftNet [45] proposes an early fusion framework that jointly encodes knowledge base facts and supporting text by constructing a heterogeneous graph, enabling effective reasoning through graph convolutional networks for open-domain question answering.
    • PullNet: PullNet [44] introduces an iterative retrieval mechanism that expands a query-specific subgraph by pulling relevant facts from both knowledge bases and text, enabling joint reasoning over heterogeneous evidence for open-domain question answering.
    • SR+NSM: SR+NSM [64] enhances multi-hop KBQA by first retrieving a question-relevant subgraph and then performing symbolic reasoning over it using Neural Symbolic Machines, improving efficiency and accuracy through constrained and focused logical inference.
    • SR+NSM+E2E: SR+NSM+E2E [64] extends SR+NSM by enabling end-to-end training that jointly optimizes subgraph retrieval and reasoning. This integration enhances model coherence and allows better alignment between retrieved subgraphs and final answer prediction.
@@
    • Flan-T5-xl: Flan-T5-xl [5] is an instruction-finetuned variant of the T5 model, trained on a diverse collection of tasks with natural language instructions. By leveraging large-scale instruction tuning, it improves zero-shot and few-shot performance across diverse NLP benchmarks.
    • Alpaca-7B: Alpaca-7B [48] is an instruction-following language model fine-tuned from LLaMA-7B using self-instruct techniques. It demonstrates strong zero-shot and few-shot performance by aligning with human instructions across various NLP tasks.
    • Llama3-8B: Llama3-8B [9] is part of the LLaMA 3 family of models, designed for improved instruction following, reasoning, and code generation. Pretrained on a high-quality corpus and fine-tuned with supervised signals, it achieves strong performance across diverse benchmarks.
    • Qwen2.5-7B: Qwen2.5-7B [49] is a 7B-parameter instruction-tuned language model developed by Alibaba, optimized for tasks such as reasoning, code generation, and dialogue. It supports multi-turn conversation and demonstrates competitive performance on standard benchmarks.
    • ChatGPT: ChatGPT [40] is a conversational AI developed by OpenAI, based on the GPT architecture. It is designed to understand natural language, engage in dialogue, answer questions, and assist with a wide range of tasks across domains.
    • ChatGPT+CoT: ChatGPT with Chain-of-Thought (CoT) [54] prompting enhances the model’s reasoning capabilities by encouraging it to generate intermediate reasoning steps before arriving at a final answer, improving performance on complex, multi-step problems.
@@
    • UniKGQA: UniKGQA [15] is a unified framework that integrates retrieval and reasoning for multi-hop question answering over knowledge graphs, combining subgraph retrieval, query decomposition, and neural reasoning in an end-to-end manner.
@@
    • DECAF: DECAF [62] is a joint framework for question answering over knowledge bases that simultaneously decodes logical forms and answers. By leveraging dual supervision, it enhances both symbolic reasoning accuracy and direct answer prediction in a unified architecture.
@@
    • KD-CoT: KD-CoT [51] is a framework that enhances the faithfulness of large language models by guiding Chain-of-Thought reasoning with external knowledge, improving accuracy in knowledge-intensive question answering tasks.
@@
    • Nutrea: Nutrea [4] proposes a neural tree search framework for context-guided multi-hop KGQA. It incrementally constructs reasoning trees by integrating question semantics and graph context, enabling efficient exploration of multi-hop paths for accurate answer prediction.
@@
    • ToG: ToG [46] is a framework that enables large language models to perform deep and responsible reasoning over knowledge graphs by combining structured graph information with iterative thinking and verification mechanisms for reliable multi-hop QA.
@@
    • RoG: RoG [28] is a framework that enhances the faithfulness and interpretability of large language model reasoning by grounding multi-hop question answering on knowledge graphs, integrating symbolic path tracking with natural language generation.
@@
    • KAPING: KAPING [1] introduces knowledge-augmented prompting by integrating structured triples into Chain-of-Thought (CoT) reasoning. It guides large language models to generate intermediate reasoning steps, enabling zero-shot multi-hop KGQA without task-specific fine-tuning.
@@
    • ReasoningLM: ReasoningLM [14] enhances pre-trained language models for KGQA by injecting subgraph structures into the input representation. It enables structural reasoning over retrieved subgraphs through a reasoning-aware encoder, improving performance on complex multi-hop queries.
@@
    • FiDeLis: FiDeLis [43] proposes a faithfulness-aware KGQA framework that enhances reasoning consistency in LLMs by aligning generated logical forms with answer predictions. It introduces fidelity constraints to reduce hallucinations and improve answer correctness.
@@
    • GNN-RAG: GNN-RAG [33] integrates graph neural networks with retrieval-augmented generation by encoding knowledge subgraphs into LLMs’ context. It enables structural reasoning over retrieved subgraphs, improving answer accuracy in KGQA through explicit graph-aware representations.
@@
    • DoG: DoG [30] is a flexible and reliable reasoning framework that enables large language models to generate and evaluate multiple reasoning paths over knowledge graphs through a debate-style process, enhancing robustness and answer faithfulness.
@@
    • DuarL: DuarL [23] is a collaborative framework that integrates GNNs and LLMs for KGQA, where GNNs capture structural semantics and LLMs perform adaptive reasoning, enabling accurate and interpretable multi-hop QA.
@@
    • DP: DP [31] is a trustworthy reasoning framework that guides large language models using prior knowledge from knowledge graphs. It iteratively verifies and refines reasoning paths to enhance faithfulness, robustness, and answer accuracy in KGQA.
@@
    • RwT: RwT [41] is a faithful KGQA framework that models multi-hop reasoning as tree-structured exploration over knowledge graphs, enabling large language models to generate interpretable reasoning paths and improve answer consistency and accuracy.
Appendix D More experimental results